Circular No. 8280

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 WC_{7} (LINEAR-CATALINA) An apparently asteroidal object was discovered by the LINEAR project

on 2003 Nov. 18 (observed on only two nights) and given the designation 2003 WC_{7} (MPS 91151). The object was discovered independently on Jan. 31 by the Catalina Sky Survey and then posted on the NEO Confirmation Page. As a result, it has been found to show cometary appearance on CCD exposures taken by J. Young (Table Mountain, 0.6-m reflector, Feb. 1.15 UT; very diffuse coma of mag 17.5 and diameter 5′′, very little central condensation, and a straight, narrow 10′′ tail in p.a. 345^{o}) and by G. J. Garradd and R. H. McNaught (Siding Spring, 1.0-m f/8 reflector, Feb. 1.46; coma diameter 3′′.5 in 2′′.5 seeing; no obvious tail visible in five co-added 40-s frames). The discovery observations follow:

Date UT α_{2000} δ_{2000} Mag. Observer
2003 Nov. 18.14403 1^{h}01^{m}07.^{s}41 −27^{o}43′24′′.9 19.6 LINEAR
2004 Jan. 31.13554 2 13 16.76 + 3 58 40.8 17.6 Catalina

New astrometry (with prediscovery Catalina data from 2003 Dec. 16), the following orbital elements, and an ephemeris appear on MPEC 2004-C03.

T = 2004 Feb. 5.7407 TT  ω = 341.^{o}0539
e = 0.680818  Ω = 89.8882 }2000.0
q = 1.654897 AU  i = 21.2195
a = 5.184810 AU n^{o} = 0.0834843 P = 11.81 years

SUPERNOVA 2004N M. Salvo and B. Schmidt, Australian National University (ANU), re-

port that a spectrogram (range 390–620 nm) of SN 2004N (cf. IAUC 8276), obtained on Jan. 29.6 UT with the ANU 2.3-m telescope (+ Double Beam Spectrograph) at Siding Spring Observatory, shows it to be a type-Ia supernova. Cross correlation with a library of type-Ia supernovae spectra indicates that SN 2004N is ≈ 1 ± 2 days before maximum light, at a redshift of z = 0.05.

COMET C/2002 T7 (LINEAR) Visual total-magnitude and coma-diameter estimates made with 50-

mm binoculars: Jan. 16.79 UT, 7.4, 11′ (J. J. Gonzalez, Asturias, Spain); 24.12, 6.9, 17′ (C. S. Morris, Fillmore, CA); 31.84, 7.5, 7′ (Gonzalez).

2004 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green